# Exploring Mutual Empowerment Between Wireless Networks and RL-based LLMs: A Survey

### v2

@@ V-B Challenges
Stability and Reliability of Network Environments. In addition to the challenges related to data and models in wireless communications, the characteristics of the network environment itself can significantly impact the effectiveness of training or deploying LLMs. The quality of data and model parameters may be compromised when transmitted over noisy wireless channels, leading to performance degradation or even loss of critical information due to communication congestion. Therefore, ensuring the robustness and efficiency of LLMs in dynamic wireless environments is essential yet remains a significant challenge. The property of the wireless environment that creates difficult problems is the swift transformations in the physical surroundings, for example, the moving object blocks the line-of-sight between the transmitter and receiver in a short time without the acknowledgment of both devices. This scenario is more likely to occur when the current wireless network prioritizes fast transmission and supports a large number of connected devices while reducing the correction protocol among them to minimize communication and computation latency. When transmitting RL-based LLMs with large sizes, the primary goal is to ensure a stable and reliable network. Otherwise, varying levels of channel noise can affect different layers in different ways, making it impossible to eliminate noise across the entire model and achieve optimal performance. Additionally, the unstable connections and fluctuating bandwidth in wireless networks can disrupt the RL training loop, resulting in incomplete learning cycles and diminished model performance. Therefore, maintaining consistent communication between the agent and the server is vital to ensure that RL-based LLMs can adapt and learn effectively in real-world applications.
@@
Fairness, Bias, and Malicious User Concerns. Due to the decentralized and heterogeneous nature of wireless networks, empowering LLMs through these networks presents significant challenges in fairness, bias, and security. First, users in a wireless network have varying levels of connectivity and resources, which can lead to fairness issues. Those with stronger connections and better resources may dominate the learning process, while users with weaker network conditions risk being underrepresented, resulting in imbalanced model optimization. Second, bias can arise if the majority of devices in the network are concentrated in specific regions, leading the model to favor certain characteristics while neglecting others. For example, if the model is predominantly trained on English data, it may struggle with resource-poor languages, generating less fluent or even inaccurate translations. Similarly, if historical hiring data exhibits a preference for male candidates in technical roles, the model may disproportionately recommend such positions to men, even when female candidates have equivalent qualifications, thereby reinforcing gender biases in technical fields. Finally, since RL-based LLMs rely on continuous interaction with users and the environment [135], they are vulnerable to malicious manipulation. Malicious users may inject poisoned data, manipulate feedback signals, or launch adversarial attacks, compromising model integrity and security. Addressing these challenges necessitates the implementation of fairness-aware learning mechanisms, bias mitigation strategies, and robust adversarial defenses to ensure that RL-based LLM frameworks in wireless networks are fair, reliable, and secure.
V-C Potential Solutions
@@
Leveraging Advanced Distributed Learning. To enhance LLMs in wireless communication systems, advanced distributed learning techniques can be employed to improve model performance and adaptability. Traditional LLMs rely on large-scale centralized data for training, necessitating compliance with data protection and privacy regulations, such as GDPR [181]. Distributed training over wireless networks, such as federated learning [65], enables LLMs to be trained across numerous local computing clusters, thereby avoiding direct centralized data storage and enhancing privacy protection capabilities. However, a risk of data leakage persists when distributed local devices collaborate to train models. For instance, during the model update process, participating devices may inadvertently reveal the statistical characteristics of their local data or model parameters, allowing attackers to infer sensitive information about certain users. To mitigate this risk, techniques such as differential privacy [197] can be implemented to ensure that individual data contributions remain confidential, while secure multi-party computation [185] can aggregate model updates without exposing any raw data. Additionally, secure aggregation protocols can help maintain data confidentiality during training.
@@
Moreover, local data sources may exhibit heterogeneity, which can negatively impact the model’s generalization ability. To address this issue, prototype-based [198, 55] domain alignment techniques can align the model with the global data distribution. Furthermore, generative artificial intelligence technologies, such as diffusion models [199, 200, 201, 202] and GANs [203, 204, 205], offer promising solutions by generating synthetic samples that simulate various data distributions, thus alleviating the effects of data heterogeneity and enhancing the model’s robustness and generalization capability.
@@
Exploring Incentive Mechanism for Distributed LLMs Development. Exploring incentive mechanisms that promote the development of distributed LLMs is critical for the advancement of the AI community, as it can effectively increase participant enthusiasm and encourage more devices and users to contribute their computing resources and data. Consequently, the design of these incentive mechanisms must be both effective and fair, taking into account the differences in computing resources and power across various regions where participants are located. In wireless networks, devices may be reluctant to participate in training or share data due to concerns about privacy, potential misuse of information, or simply a lack of perceived benefits. Therefore, future technical designs should provide an open and transparent framework for the use of data and computing resources, while also offering appropriate rewards to users who contribute, such as free access to LLM API resources. By establishing a balanced incentive structure, we can foster a collaborative environment that promotes the development of distributed LLMs while ensuring user participation and data protection.
@@
Next-generation Communication Technologies. The development of LLMs enabled by wireless network systems presents challenges such as network bandwidth limitations, which can hinder model training and updates. Therefore, next-generation communication technologies such as 6G [206], semantic communication [207], and integrated sensing and communication (ISAC) [23] offer promising solutions. 6G technology is designed to provide ultra-reliable low-latency communication and significantly increased data rates, facilitating the seamless transmission of LLMs updates and real-time interactions. Semantic communication improves efficiency by focusing on the transmission of meaningful information rather than raw data, thereby significantly reducing bandwidth requirements and latency. This is especially beneficial for LLMs, as it enables faster and more efficient data exchange. Additionally, ISAC integrates communication with environmental sensing, allowing for optimized resource utilization and enhanced situational awareness. By leveraging these advanced communication technologies, we can establish a more robust framework for distributed LLMs, ensuring effective model performance while addressing critical issues related to data privacy and computational demands.
@@
Figure 13: Illustration of applications demonstrating the mutual empowerment between wireless communication systems and LLMs.
@@
Trustworthy and Robust Model Design. In order to effectively enable LLMs over wireless networks, it is essential to design trustworthy and robust models that address key challenges such as unfairness, bias, and unreliable environments. Trustworthy AI [208] emphasizes the reliability, transparency, and ethical considerations of AI systems, ensuring that users can trust the decisions made by these systems. Training models based on data from a specific region can lead to biased decisions, reinforcing stereotypes that associate engineers with men and nurses with women. While distributed model training can help mitigate this risk, the black-box nature of model decisions often leaves users unaware of the underlying processes. This lack of understanding can foster doubts about the model’s fairness, leading to a lack of trust in its outputs and potentially exacerbating bias and discrimination against certain groups. Therefore, the development of these models should incorporate several core elements: transparency, which enables users to understand the reasoning behind AI decisions; fairness, which seeks to minimize bias and discrimination; privacy protection, ensuring user data remains secure; and traceability, allowing the decision-making process to be tracked.
@@
In addition, robust AI [209] focuses on the stability and adaptability of AI systems under various conditions, particularly in the face of uncertainty or interference. This encompasses maintaining performance despite disturbances, adapting to environmental changes and new data distributions, and demonstrating fault tolerance to manage errors or anomalies without significant failures. For instance, to mitigate the impact of adversarial attacks [3], adversarial samples can be intentionally introduced during the training phase to reduce the likelihood of successful attacks. By creating models that users can trust, that perform well across diverse environments, and that adhere to ethical standards, we can enhance user trust and participation in LLM applications over wireless networks.
@@
Green Model Design The development of LLMs demands substantial hardware and energy resources, significantly increasing their carbon footprint. Therefore, designing green models is essential for building sustainable AI systems. To minimize environmental impact, strategies such as quantization [175], model sharding [210], knowledge distillation [3], and efficient architectural designs can be employed. Quantization reduces the storage and computational requirements of a model by lowering the bit precision of weights and activations, thereby decreasing energy consumption and carbon emissions. Model sharding partitions large models into smaller, more manageable segments, enabling distributed devices to collaboratively train different model components. This not only alleviates the computational burden on individual devices but also enhances training efficiency by allowing each device to focus on a specific shard. Knowledge distillation further optimizes resource usage by transferring knowledge from a larger, complex teacher model to a smaller, lightweight student model, ensuring efficient inference with reduced computational costs.
@@
In addition, efficient architectural design plays a critical role in reducing carbon emissions. For instance, in long-text modeling, traditional full attention mechanisms exhibit quadratic complexity with respect to sequence length, leading to high memory and computation costs. Recent advancements in native sparse attention [211] have optimized conventional full attention, reducing memory access to just 1/11.6 of its original requirement without compromising performance. By integrating these technologies into wireless network systems, we can advance LLM development while ensuring a more energy-efficient and sustainable approach, allowing these models to operate effectively in diverse environments with minimal environmental impact. An overview of potential solutions is illustrated in Fig. 12.
VI Applications
@@
In this section, we explore four applications as illustrated in Fig. 13, including intelligent ecosystems, autonomous vehicular networks, smart manufacturing and Industry 4.0, and personalized education. These examples highlight the mutual empowerment of wireless networking systems and large language models (LLMs). While multiple other applications may exist, these four serve as representative cases to facilitate discussion.
VI-A Intelligent Ecosystems
@@
The integration of wireless network systems with LLMs is revolutionizing smart ecosystems across various domains. In smart cities, LLMs facilitate real-time data processing from IoT sensors to optimize traffic management, energy distribution, and public safety, while wireless networks ensure seamless communication between interconnected devices. This continuous stream of data can further enhance LLM performance. Similarly, in smart healthcare, AI-driven diagnostics and automated patient monitoring leverage wireless connectivity to enable remote medical services and real-time consultations. In smart homes, LLM-powered virtual assistants, such as Apple Home and LLM Siri, can improve voice recognition, personalized automation, and security system integration, creating a more intuitive living experience. Moreover, in smart agriculture, the synergy between LLM-driven analytics and wireless sensor networks fosters precision agriculture by optimizing irrigation, pest control, and crop yield predictions, ultimately boosting productivity. This mutual empowerment also extends to wearable technology, which can provide real-time language translation, biometric monitoring, and personalized user experiences, with wireless connectivity ensuring seamless data exchange and continuous updates. By combining the intelligence of LLMs with the adaptability of wireless infrastructure, these smart ecosystems would become more autonomous, responsive, and mission-critical, driving advancements in automation, decision-making, and operational efficiency across diverse applications.
VI-B Autonomous Vehicular Networks
@@
Developing an autonomous vehicular network involves a multifaceted development that encompasses numerous challenges. Advanced sensing technologies, such as LIDAR, radar, and cameras, are crucial for accurate perception and decision-making, requiring sophisticated AI algorithms for effective operation. Real-time response is essential for the effective operation of these systems. Integrating LLMs into wireless communication can mitigate the burden of processing data by employing semantic communication and contextual understanding. Moreover, LLMs play a vital role in optimizing network traffic in both wireless communication and vehicular routing. By selecting the most efficient routes for data transmission, LLMs can enhance real-time response capabilities while providing valuable information for AI to control the vehicle.
@@
For example, LLMs can analyze and interpret complex data from various sensors and communication channels, enabling smarter decision-making and enhanced situational awareness. Additionally, wireless networks facilitate dynamic communication between vehicles and infrastructure, providing real-time updates on traffic conditions, potential hazards, and optimal routes. This synergy is expected to lead to a safer and more efficient network of autonomous vehicles that can adapt to changing environments and user needs. Recent research introduces DriveGPT4 [212], an interpretable end-to-end autonomous driving system that extends LLMs into the multimodal domain to process road scene data. DriveGPT4 has demonstrated promising performance in predicting low-level vehicle control signals, including steering, acceleration, braking, and gear shifting.
VI-C Industry 4.0
@@
Digital twins have become essential in smart manufacturing and Industry 4.0, changing how industries work. One new trend is using real-time AI for industrial automation, which involves creating an AI gym where robots and AI agents can be trained and tested in complex environments such as storage facilities [213]. This requires many techniques, including LLMs, for control, decision-making, sensing, and communication, as well as wireless connections for smooth interactions between agents and servers.
@@
For example, in storage management, an AI robot needs to load and organize products in specific locations. The path from loading to unloading has many variables that can affect the efficiency of this action. To handle this, the AI agent collects data from sensors, cameras, microphones, and other sources, sending this information to the server for processing. Any delays in sending data or commands can cause financial losses and other problems. Therefore, it’s crucial to have a strong system that can make quick and accurate decisions for these applications using LLM and wireless networks.
VI-D Personalized Education

### v3

@@ 4.2.3. Stability and Reliability of Network Environments
In addition to the challenges related to data and models in wireless communications, the characteristics of the network environment itself can significantly impact the effectiveness of training or deploying LLMs. The quality of data and model parameters may be compromised when transmitted over noisy wireless channels, leading to performance degradation or even loss of critical information due to communication congestion (Javaid et al., 2021). Therefore, ensuring the robustness and efficiency of LLMs in dynamic wireless environments is essential yet remains a significant challenge. The property of the wireless environment that creates difficult problems is the swift transformations in the physical surroundings, for example, the moving object blocks the line-of-sight between the transmitter and receiver in a short time without the acknowledgment of both devices (Bensky, 2019). This scenario is more likely to occur when the current wireless network prioritizes fast transmission and supports a large number of connected devices while reducing the correction protocol among them to minimize communication and computation latency. When transmitting RL-based LLMs with large sizes, the primary goal is to ensure a stable and reliable network. Otherwise, varying levels of channel noise can affect different layers in different ways, making it impossible to eliminate noise across the entire model and achieve optimal performance. Additionally, the unstable connections and fluctuating bandwidth in wireless networks can disrupt the RL training loop, resulting in incomplete learning cycles and diminished model performance (Chen et al., 2020). Therefore, maintaining consistent communication between the agent and the server is vital to ensure that RL-based LLMs can adapt and learn effectively in real-world applications.
4.2.4. Fairness, Bias, and Malicious User Concerns
@@
Due to the decentralized and heterogeneous nature of wireless networks, empowering LLMs through these networks presents significant challenges in fairness, bias, and security. First, users in a wireless network have varying levels of connectivity and resources, which can lead to fairness issues (Wu et al., 2025). Those with stronger connections and better resources may dominate the learning process, while users with weaker network conditions risk being underrepresented, resulting in imbalanced model optimization. Second, bias can arise if the majority of devices in the network are concentrated in specific regions, leading the model to favor certain characteristics while neglecting others (Javed et al., 2025). For example, if the model is predominantly trained on English data, it may struggle with resource-poor languages, generating less fluent or even inaccurate translations. Similarly, if historical hiring data exhibits a preference for male candidates in technical roles, the model may disproportionately recommend such positions to men, even when female candidates have equivalent qualifications, thereby reinforcing gender biases in technical fields (Mujtaba and Mahapatra, 2024). Finally, since RL-based LLMs rely on continuous interaction with users and the environment, they are vulnerable to malicious manipulation (Hurst et al., 2024). Malicious users may inject poisoned data, manipulate feedback signals, or launch adversarial attacks (Qiao et al., 2025b), compromising model integrity and security. Addressing these challenges necessitates the implementation of fairness-aware learning mechanisms, bias mitigation strategies, and robust adversarial defenses to ensure that RL-based LLM frameworks in wireless networks are fair, reliable, and secure.
4.3. Potential Solutions
Figure 11. Potential solutions for empowering LLMs with wireless communication systems.
4.3.1. Leveraging Advanced Distributed Learning
@@
To enhance LLMs in wireless communication systems, advanced distributed learning techniques can be employed to improve model performance and adaptability. Traditional LLMs rely on large-scale centralized data for training, necessitating compliance with data protection and privacy regulations, such as GDPR (Voigt and Von dem Bussche, 2017). Distributed training over wireless networks, such as federated learning (McMahan et al., 2017), enables LLMs to be trained across numerous local computing clusters, thereby avoiding direct, centralized data storage and enhancing privacy protection capabilities. However, a risk of data leakage persists when distributed local devices collaborate to train models. For instance, during the model update process, participating devices may inadvertently reveal the statistical characteristics of their local data or model parameters, allowing attackers to infer sensitive information about certain users. To mitigate this risk, techniques such as differential privacy (Abadi et al., 2016) can be implemented to ensure that individual data contributions remain confidential, while secure multi-party computation (Zhou et al., 2024b) can aggregate model updates without exposing any raw data. Additionally, secure aggregation protocols can help maintain data confidentiality during training.
@@
Moreover, local data sources may exhibit heterogeneity, which can negatively impact the model’s generalization ability. To address this issue, prototype-based (Le et al., 2025; Qiao et al., 2025b) domain alignment techniques can align the model with the global data distribution. Furthermore, generative artificial intelligence technologies, such as diffusion models (Croitoru et al., 2023; Sohl-Dickstein et al., 2015; Nguyen et al., 2023a; Ho et al., 2020) and GANs (Xia et al., 2022; Goodfellow et al., 2014; Mao et al., 2017a), offer promising solutions by generating synthetic samples that simulate various data distributions, thus alleviating the effects of data heterogeneity and enhancing the model’s robustness and generalization capability.
4.3.2. Exploring Incentive Mechanism for Distributed LLMs Development
@@
Exploring incentive mechanisms that promote the development of distributed LLMs is critical for the advancement of the AI community, as it can effectively increase participant enthusiasm and encourage more devices and users to contribute their computing resources and data (Tran et al., 2025). Consequently, the design of these incentive mechanisms must be both effective and fair, taking into account the differences in computing resources and power across various regions where participants are located. In wireless networks, devices may be reluctant to participate in training or share data due to concerns about privacy, potential misuse of information, or simply a lack of perceived benefits (Chen et al., 2024d). Therefore, future technical designs should provide an open and transparent framework for the use of data and computing resources while also offering appropriate rewards to users who contribute, such as free access to LLM API resources. By establishing a balanced incentive structure, we can foster a collaborative environment that promotes the development of distributed LLMs while ensuring user participation and data protection.
4.3.3. Next-generation Communication Technologies
@@
The development of LLMs enabled by wireless network systems presents challenges such as network bandwidth limitations, which can hinder model training and updates. Therefore, next-generation communication technologies such as 6G (Jiang et al., 2021), semantic communication (Yang et al., 2022), and integrated sensing and communication (ISAC) (Adhikary et al., 2023b, 2022) offer promising solutions. 6G technology is designed to provide ultra-reliable low-latency communication and significantly increased data rates, facilitating the seamless transmission of LLMs updates and real-time interactions. Semantic communication improves efficiency by focusing on the transmission of meaningful information rather than raw data, thereby significantly reducing bandwidth requirements and latency. This is especially beneficial for LLMs, as it enables faster and more efficient data exchange. Additionally, ISAC integrates communication with environmental sensing, allowing for optimized resource utilization and enhanced situational awareness. By leveraging these advanced communication technologies, we can establish a more robust framework for distributed LLMs, ensuring effective model performance while addressing critical issues related to data privacy and computational demands.
@@
Figure 12. Illustration of applications demonstrating the mutual empowerment between wireless communication systems and LLMs.
4.3.4. Trustworthy and Robust Model Design
@@
In order to effectively enable LLMs over wireless networks, it is essential to design trustworthy and robust models that address key challenges such as unfairness, bias, and unreliable environments. Trustworthy AI (Li et al., 2023) emphasizes the reliability, transparency, and ethical considerations of AI systems, ensuring that users can trust the decisions made by these systems. Training models based on data from a specific region can lead to biased decisions, reinforcing stereotypes that associate engineers with men and nurses with women. While distributed model training can help mitigate this risk, the black-box nature of model decisions often leaves users unaware of the underlying processes. This lack of understanding can foster doubts about the model’s fairness, leading to a lack of trust in its outputs and potentially exacerbating bias and discrimination against certain groups. Therefore, the development of these models should incorporate several core elements: transparency, which enables users to understand the reasoning behind AI decisions; fairness, which seeks to minimize bias and discrimination; privacy protection, ensuring user data remains secure; and traceability, allowing the decision-making process to be tracked.
@@
In addition, robust AI (Tocchetti et al., 2022) focuses on the stability and adaptability of AI systems under various conditions, particularly in the face of uncertainty or interference. This encompasses maintaining performance despite disturbances, adapting to environmental changes and new data distributions, and demonstrating fault tolerance to manage errors or anomalies without significant failures. For instance, to mitigate the impact of adversarial attacks (Qiao et al., 2025a), adversarial samples can be intentionally introduced during the training phase to reduce the likelihood of successful attacks. By creating models that users can trust, that perform well across diverse environments, and that adhere to ethical standards, we can enhance user trust and participation in LLM applications over wireless networks.
4.3.5. Green Model Design
@@
The development of LLMs demands substantial hardware and energy resources, significantly increasing their carbon footprint. Therefore, designing green models is essential for building sustainable AI systems. To minimize environmental impact, strategies such as quantization (Shao et al., 2024a), model sharding (Quan et al., 2024), knowledge distillation (Qiao et al., 2025a), and efficient architectural designs can be employed. Quantization reduces the storage and computational requirements of a model by lowering the bit precision of weights and activations, thereby decreasing energy consumption and carbon emissions. Model sharding partitions large models into smaller, more manageable segments, enabling distributed devices to train different model components collaboratively. This not only alleviates the computational burden on individual devices but also enhances training efficiency by allowing each device to focus on a specific shard. Knowledge distillation further optimizes resource usage by transferring knowledge from a larger, complex teacher model to a smaller, lightweight student model, ensuring efficient inference with reduced computational costs.
@@
In addition, efficient architectural design plays a critical role in reducing carbon emissions. For instance, in long-text modeling, traditional full attention mechanisms exhibit quadratic complexity with respect to sequence length, leading to high memory and computation costs. Recent advancements in native sparse attention (Yuan et al., 2025) have optimized conventional full attention, reducing memory access to just 1/11.6 of its original requirement without compromising performance. By integrating these technologies into wireless network systems, we can advance LLM development while ensuring a more energy-efficient and sustainable approach, allowing these models to operate effectively in diverse environments with minimal environmental impact. An overview of potential solutions is illustrated in Fig. 11.
5. Applications
@@
In this section, we explore four applications as illustrated in Fig. 12, including intelligent ecosystems, autonomous vehicular networks, smart manufacturing and Industry 4.0, and personalized education. These examples highlight the mutual empowerment of wireless networking systems and LLMs. While multiple other applications may exist, these four serve as representative cases to facilitate discussion.
5.1. Intelligent Ecosystems
@@
The integration of wireless network systems with LLMs is revolutionizing smart ecosystems across various domains. In smart cities (Kalyuzhnaya et al., 2025), LLMs facilitate real-time data processing from IoT sensors to optimize traffic management, energy distribution, and public safety, while wireless networks ensure seamless communication between interconnected devices. This continuous stream of data can further enhance LLM performance. Similarly, in smart healthcare (Lai and Lai, 2022), AI-driven diagnostics and automated patient monitoring leverage wireless connectivity to enable remote medical services and real-time consultations. In smart homes (Rivkin et al., 2024), LLM-powered virtual assistants, such as Apple Home and LLM Siri, can improve voice recognition, personalized automation, and security system integration, creating a more intuitive living experience. Moreover, in smart agriculture (Li et al., 2024b), the synergy between LLM-driven analytics and wireless sensor networks fosters precision agriculture by optimizing irrigation, pest control, and crop yield predictions, ultimately boosting productivity. This mutual empowerment also extends to wearable technology, which can provide real-time language translation, biometric monitoring, and personalized user experiences, with wireless connectivity ensuring seamless data exchange and continuous updates. By combining the intelligence of LLMs with the adaptability of wireless infrastructure, these smart ecosystems would become more autonomous, responsive, and mission-critical, driving advancements in automation, decision-making, and operational efficiency across diverse applications.
5.2. Autonomous Vehicular Networks
@@
Developing an autonomous vehicular network involves a multifaceted development that encompasses numerous challenges. Advanced sensing technologies, such as LIDAR, radar, and cameras, are crucial for accurate perception and decision-making, requiring sophisticated AI algorithms for effective operation (Thakur and Mishra, 2024). Real-time response is essential for the effective operation of these systems. Integrating LLMs into wireless communication can mitigate the burden of processing data by employing semantic communication and contextual understanding. Moreover, LLMs play a vital role in optimizing network traffic in both wireless communication and vehicular routing (Hu et al., 2025). By selecting the most efficient routes for data transmission, LLMs can enhance real-time response capabilities while providing valuable information for AI to control the vehicle.
@@
For example, LLMs can analyze and interpret complex data from various sensors and communication channels, enabling smarter decision-making and enhanced situational awareness (Hu et al., 2025). Additionally, wireless networks facilitate dynamic communication between vehicles and infrastructure, providing real-time updates on traffic conditions, potential hazards, and optimal routes. This synergy is expected to lead to a safer and more efficient network of autonomous vehicles that can adapt to changing environments and user needs. Recent research introduces DriveGPT4 (Xu et al., 2024c), an interpretable end-to-end autonomous driving system that extends LLMs into the multimodal domain to process road scene data. DriveGPT4 has demonstrated promising performance in predicting low-level vehicle control signals, including steering, acceleration, braking, and gear shifting.
5.3. Industry 4.0
@@
Digital twins have become essential in smart manufacturing and Industry 4.0, changing how industries work (nvi, [n. d.]). One new trend is using real-time AI for industrial automation, which involves creating an AI gym where robots and AI agents can be trained and tested in complex environments such as storage facilities (nvi, [n. d.]). This requires many techniques, including LLMs, for control, decision-making, sensing, and communication, as well as wireless connections for smooth interactions between agents and servers.
@@
For example, in storage management, an AI robot needs to load and organize products in specific locations (Murphy, 2019). The path from loading to unloading has many variables that can affect the efficiency of this action. To handle this, the AI agent collects data from sensors, cameras, microphones, and other sources, sending this information to the server for processing. Any delays in sending data or commands can cause financial losses and other problems. Therefore, it’s crucial to have a strong system that can make quick and accurate decisions for these applications using LLM and wireless networks.
5.4. Personalized Education